the side of a polygon are 6, 10, 20, and 24 inches long. if the shortest side is reduced by 2 inches, what reductions should be made in the other side to obtain a similar polygon?
Answer by jim_thompson5910(35256)

The shortest side of the original polygon is 6

"the shortest side is reduced by 2 inches" so 6 turns into 6-2 = 4

Divide 4 by 6 to get 4/6 = 2/3. So this means that we must multiply every dimension by 2/3 to get a smaller similar polygon. So for example, the side 10 turns into 10*2/3 = 20/3 = 6 & 2/3. The same applies for the other dimensions.